OpenStudy (anonymous):

Stuck on this integration using 'u' substitution.. integrate (e^x)/(sqrt(4+e^2x))

OpenStudy (anonymous):

substitute e^x=t

OpenStudy (anonymous):

e^2x is also (e^x)^2?

OpenStudy (anonymous):

\[\int\limits_{}^{}e ^{x}dx/\sqrt{4+(e ^{x})^{2}}\] put e^x=t derivate both sides wrt to x (e^x) dx=dt \[\int\limits_{}^{}dt/\sqrt{(2)^{2}+t ^{2}}\]

OpenStudy (anonymous):

now use the formula : \[\int\limits_{}^{}dt/\sqrt{a ^{2}+t ^{2}}=(1/a)\log \left| (a+t)/(a-t) \right|\]

OpenStudy (anonymous):

in final answer dont forget to put t=e^x back

OpenStudy (anonymous):

Got it, thanks. It was e^2x throwing me off, forgot you could use (e^x)^2...thanks again

OpenStudy (anonymous):

ur welcome